OpenStudy (anonymous):

Which of the following equations has only one solution? x2- 8x+ 16 = 0 x(x- 1) = 8 x2= 16

OpenStudy (anonymous):

x2- 8x+ 16 = 0

OpenStudy (mathmate):

All of them have two solutions. One of them has two solutions that have the same value. Factorize the first one: (x-4)(x-4)=0 and you'll find which one has "one" solution.

OpenStudy (waleed_imtiaz):

yes...... Its true...... Have the same value if we factorize.... So x=4 .... will be the answer of the equation x2- 8x+ 16 = 0